# Module schemes in invariant theory

### v1

@@ Introduction
Let $\rho$ be another linear representation of $G$ in $\operatorname{Gl}_{m}(K)$ with the corresponding linear action on a vector space $V\cong K^{m}$ or on the affine space ${\mathbb{A}}^{m}$. This defines the module of covariants $(S\otimes_{K}V)^{G}$ ($G$ acting on both tensor components) which is a maximal Cohen-Macaulay module over $R$. If $G$ is a small group, meaning that $\beta(G)\subseteq\operatorname{Gl}_{d}(K)$ does not contain any (pseudo-)reflection (a linear mapping $\neq\operatorname{Id}$ that fixes a hyperplane), then we have the following correspondence (see [30, Corollary 5.20]).
@@
Section 20 deals with relations between reflections of the basic representation and reflections of a product representations. This is helpful in Section 21 to understand (for $K={\mathbb{C}}$) where the singularities of a quotient scheme $Z$ are located and how they are related to the singularities of $X$. In Section 22, we compute the fundamental group of the regular locus of $Z_{\rho}$ in terms of the acting group $G$ and the reflection subgroup of the product representation (Theorem 22.10).
@@
I thank Marc Chardin, Hailong Dao, David Eisenbud, Eleonore Faber, Robin Hartshorne, Craig Huneke, Colin Ingalls, Ines Melzer, Mandira Mondal, Stefan Schröer, Bernd Sturmfels and Bernd Ulrich for useful discussions, comments and references.
@@ 1. The setting
The vertical morphisms are quotient morphisms for group actions, and the horizontal morphisms are bundle-like morphisms (which has to be made precise for the lower morphism, see Section 12). These objects $Z_{\rho}\rightarrow X$ are the main objects of this paper. More precisely, for $G$, $Y$ and $X$ fixed, we want to understand how the $Z_{\rho}$ reflect properties of the (varying) representations $\rho$. This is also the reason why, in the second component, we stick to linear actions on a polynomial ring. The schemes $Z_{\rho}$ are always normal, as this is true for all quotient schemes of a normal scheme under a finite group, and the invariant algebra is always an ${\mathbb{N}}$-graded (not standard-graded) Cohen-Macaulay (in the nonmodular case, i.e., if the group order is not 00 in $K$, see [29, Proposition 13]) and of finite type over $K$ if $Y$ is. If the representation $\rho$ is trivial, then $Z_{\rho}=X\times{\mathbb{A}}^{m}$ is the trivial vector bundle over $X$ ($m=0$ yields the zero-bundle $X\rightarrow X$).
@@ 2. Some lemmata
The following is basically proven in [33, Theorem 1.1] for reductive groups in char 00 (see [33, Amplification 1.3] for the nonmodular case in positive characteristic) to give a universal geometric quotient, but we need base change on the quotient itself (for finite groups).
@@
In the free case according to [33, Proposition 0.9], $Y\rightarrow X$ is faithfully flat and $G\times Y\rightarrow Y\times_{X}Y$ is a $G$-equivariant isomorphism, where the action on the left is only on the first component. Such actions are called principal fiber bundles. As we consider only group schemes $G$ coming from a usual group (so that the corresponding group scheme is reduced), the morphism $Y\rightarrow X$ in the free case (and if everything is essentially of finite type) is also étale.
@@
where the lower row represents the setting of faithfully flat descent, see [37, Section VIII]. When the action is free, we have vertical isomorphisms everywhere, see [6, Example 6.2.B], and then the data above $Y$ (a module, an algebra, an affine scheme) with a group action are compatible with the group action if and only if they are descent data.
@@ 5. Invariant modules and invariant algebras
We have $S\otimes_{R}S\cong K^{|G|}\otimes S\cong S^{|G|}$ by [33, Proposition 0.9] and $R\rightarrow S$ is faithfully flat. We look at the diagram
@@
We now use descent properties for faithfully flat base change. By [24, Proposition 2.7.1 (viii)], it follows that the original ring homomorphism is an isomorphism, and by [24, Proposition 2.5.2] it follows that $M$ is locally free. ∎
@@
The previous result can also be deduced from faithfully flat descent for affine schemes (see [37, Théorème VIII.2.1]). To see this, one needs the mentioned result from GIT that $R\subseteq S$ is flat and that the compatible free group action corresponds to descent data, see [6, Example 6.2.B].
@@ 7. Ramification
Let $H\neq G$ be the kernel of the representation $\rho$. We mod out everywhere $H$, which gives the same situation back, with the extra property that $G/H$ acts faithfully on ${\mathbb{A}}^{m}$. Note that $Y/H$ is also nonreduced, otherwise, it would already be $\operatorname{Spec}K$. So we may assume that $G$ acts faithfully on ${\mathbb{A}}^{m}$. Let $V\subseteq{\mathbb{A}}^{m}$ be a nonempty open subset where the action is free, which exists by Lemma 2.3. We consider the action of $G$ on $Y\times V$. This action is also free by Lemma 2.4, so $Y\times V\rightarrow(Y\times V)/G$ is étale by Remark 2.2. Since $Y\times V$ is not reduced, $(Y\times V)/G$ is also not reduced by étale descent, see [42, Lemma 10.163.7], hence, $(Y\times{\mathbb{A}}^{m})/G$ is not reduced because it contains $(Y\times V)/G$ as an open subset. ∎
@@ 9. Regular representation
where the index set runs through a system $C$ of all irreducible representations, see [40, Corollary 1 in I.2.4]. $G$ acts on the polynomial ring $S[K^{|G|}]=\operatorname{Sym}(S^{|G|})$ with invariant algebra $(S[K^{|G|}])^{G}$. Here, $g$ sends the variable $W_{h}$ to $W_{hg}$.
@@ 10. Module schemes
Let $X$ denote a scheme. A well-known correspondence is that between (geometric) vector bundles over $X$ and locally free ${\mathcal{O}}_{X}$-modules on $X$. A vector bundle is a scheme $Z\rightarrow X$ that looks locally like $X\times{\mathbb{A}}^{r}\rightarrow X$ and has the property that the transition mappings are linear. A locally free ${\mathcal{O}}_{X}$-module looks locally like ${\mathcal{O}}_{X}^{r}$. The correspondence assigns to $Z$ the dual sheaf of the sheaf of sections and to a locally free sheaf $\mathcal{F}$ the (relative) spectrum of the symmetric algebra $\operatorname{Sym}({\mathcal{F}})=\oplus_{n\in{\mathbb{N}}}\operatorname{Sym}^{n}({\mathcal{F}})$, [26, Exercise II.5.18]. This last definition can be applied for any quasicoherent ${\mathcal{O}}_{X}$-module $\mathcal{F}$, see [23, Definition 1.7.8], and [23, Section 1.3] for the relative spectrum of a quasicoherent algebra.
@@
This formulation requires that we consider ${\mathbb{A}}^{1}_{X}$ with the natural addition and multiplication as a ring scheme (this term appears also in [34, Section 26], the affine line in the only ring scheme we are dealing with). If $\alpha:Z\times_{X}Z\rightarrow Z$ denotes the addition of the group scheme, then this definition means, to give some examples, that the diagram
@@
The concept of a module scheme is referred to briefly in [22, Chapitre 0.8] (in particular 0.8.2.3), Grothendieck writes “L’exemple des groupes a été traité avec assez de détails, mais par la suite nous laisserons généralement au lecteur le soin de développer des considérations analogues dans les exemples de structures algébriques que nous rencontrerons” (see [22, 8.2.8]). The concept is also mentioned in [14, 4.3.3], but a thorough study is missing. Related work, but only over a field, has been done by Greenberg ([20], [21], where module variety over a (noncommutative) ring variety is used) and, in the context of graded Hop algebras, by Milnor and Moore ([32], see also Remark 11.4 below).
The following lemma is stated in [23, 1.7.13] (see also [25, Section 9.4]. Grothendieck sets $V({\mathcal{E}})=\operatorname{Spec}(\operatorname{Sym}{\mathcal{E}})$ and calls this, quite inappropriately, fibré vectoriel ([14] uses fibration vectorielle instead), and writes “Nous interpréterons ces faits en disant que $S[T]$ est un $S$-schéma d’anneaux et que $V({\mathcal{E}})$ est un $S$-schéma de modules sur le S-schéma d’anneaux (cf. chap. 0, § 8)”.
@@ 11. Module schemes and graded Hopf algebras
Positively graded Hopf-algebras $A$ with $A_{0}=K$ a field (called connected) appear in the work of Milnor and Moore. If $K$ has characteristic 00, they prove results (the theorem of Milnor-Moore) that imply (under commutative and cocommutative assumption) that $A$ is a polynomial algebra over $K$, generated by homogeneous elements, see [32, Corollary 4.18] and [12, Theorem 3.8.3]. Translated into Grothendiecks terminology, this means via Lemma 11.1 and Lemma 11.3 that a module scheme over the spectrum of a field of characteristic zero is a vector space (the spectrum of the symmetric algebra of a vector space).
@@ 12. Module schemes up to normalization
The data that constitute a module scheme up to normalization are determined by the data on the restriction to $U$, by [25, Proposition 7.2.2]. The only relevant question is whether these data can be extended to some $\widetilde{Z\times_{X}Z}$. Since the properties are expressed by the commutativity of certain diagrams, i.e., identities between morphisms, these identities only need to be checked on a dense open subset.
@@ 13. Linear actions on module schemes
(1) is clear. (2) and (3) follow from Lemma 12.5. (4) is an application of [25, Proposition 7.2.2]. For $U$ sufficiently small, we get module schemes. ∎
@@
A symmetric algebra $\operatorname{Spec}\operatorname{Sym}M\rightarrow\operatorname{Spec}R$ for a finitely generated $R$-module $M$ comes naturally together with a linear injective mapping $\operatorname{Spec}\operatorname{Sym}M\rightarrow\operatorname{Spec}R\times{\mathbb{A}}^{\nu}$ above $\operatorname{Spec}R$, defined by a surjection $R^{\nu}\rightarrow M$, where $\nu$ is the number of generators of the module. If $M$ is locally free above $U$, then we get, by [25, Théorem 9.7.4], on $U$ a mapping $\vartheta:U\rightarrow\operatorname{Gr}(\nu+r,r)$ into the Grassmann variety, where $r$ denotes the rank of $M$, and where the pull-back of the universal bundle on $\operatorname{Gr}(\nu+r,r)$ gives back the embedding of the bundle over $U$. The higherdimensional fibers of $\operatorname{Spec}\operatorname{Sym}M\rightarrow\operatorname{Spec}R$ outside of $U$ are reflected by the property that the closure of the image of $\vartheta$ degenerates to a higher dimension. For a module scheme $\operatorname{Spec}A\rightarrow\operatorname{Spec}R$ up to normalization, realizing the same module $M$ above $U$, the embedding $\operatorname{Spec}A\subseteq\operatorname{Spec}R\times{\mathbb{A}}^{\mu}$ (where $\mu$ denotes the number of $R$-algebra generators of $A$) is not linear in general (not even above $U$), as Example 3.10 shows. Hence, there is also no interpretation of $\operatorname{Spec}A$ using the Grassmannian.
@@ 16. Similar modifications
In the construction of abelian varieties by A. Weil, one first has only a rational composition $V\times V\cdots\rightarrow V$, which must then be extended to a proper variety (see [47, Théorème 15], [14, Exposé XVIII]), and a similar construction occurs if one wants to spread an abelian variety defined over a quotient field $Q(R)$ to an abelian scheme over a $R$, see [6, Section 4.3].

### v2

@@ Introduction
Let $\rho$ be another linear representation of $G$ in $\operatorname{Gl}_{m}(K)$ with the corresponding linear action on a vector space $V\cong K^{m}$ or on the affine space ${\mathbb{A}}^{m}$. This defines the module of covariants $(S\otimes_{K}V)^{G}$ ($G$ acting on both tensor components) which is a maximal Cohen-Macaulay module over $R$. If $G$ is a small group, meaning that $\beta(G)\subseteq\operatorname{Gl}_{d}(K)$ does not contain any (pseudo-)reflection (a linear mapping $\neq\operatorname{Id}$ that fixes a hyperplane) and nonmodular, meaning that its order is not a multiple of the characteristic of the field, then we have the following correspondence (see [32, Corollary 5.20]).
@@
Section 20 deals with relations between reflections of the basic representation and reflections of a product representations. This is helpful in Section 21 to understand where the singularities of a quotient scheme $Z$ are located and how they are related to the singularities of $X$. In Section 22, we compute for $K={\mathbb{C}}$ the fundamental group of the regular locus of $Z_{\rho}$ in terms of the acting group $G$ and the reflection subgroup of the product representation (Theorem 22.10).
@@
I thank Janusz Adamus, Severin Barmeier, Marc Chardin, Hailong Dao, David Eisenbud, Eleonore Faber, Sudarshan Gurjar, Robin Hartshorne, Mitsuyasu Hashimoto, Craig Huneke, Najib Idrissi-Kaitouni, Colin Ingalls, Christian Liedtke, Ines Melzer, Claudia Miller, Rosa Miró-Roig, Mandira Mondal, Julia Pevtsova, Peter Schenzel, Stefan Schröer, Anurag Singh, Ilya Smirnov, Bernd Sturmfels, Peter Symonds, Shunsuke Takagi, Vijaylaxmi Trivedi, Bernd Ulrich and Keiichi Watanabe for useful discussions, comments and references.
@@ 1. The setting
The vertical morphisms are quotient morphisms for group actions, and the horizontal morphisms are bundle-like morphisms (which has to be made precise for the lower morphism, see Section 12). These objects $Z_{\rho}\rightarrow X$ are the main objects of this paper. More precisely, for $G$, $Y$ and $X$ fixed, we want to understand how the $Z_{\rho}$ reflect properties of the (varying) representations $\rho$. This is also the reason why, in the second component, we stick to linear actions on a polynomial ring. The schemes $Z_{\rho}$ are always normal, as this is true for all quotient schemes of a normal scheme under a finite group, and the invariant algebra is always an ${\mathbb{N}}$-graded (not standard-graded) Cohen-Macaulay (in the nonmodular case, i.e., if the group order is not 00 in $K$, see [31, Proposition 13]) and of finite type over $K$ if $Y$ is. If the representation $\rho$ is trivial, then $Z_{\rho}=X\times{\mathbb{A}}^{m}$ is the trivial vector bundle over $X$ ($m=0$ yields the zero-bundle $X\rightarrow X$).
@@ 2. Some lemmata
The following is basically proven in [35, Theorem 1.1] for reductive groups in char 00 (see [35, Amplification 1.3] for the nonmodular case in positive characteristic) to give a universal geometric quotient, but we need base change on the quotient itself (for finite groups).
@@
In the free case according to [35, Proposition 0.9], $Y\rightarrow X$ is faithfully flat and $G\times Y\rightarrow Y\times_{X}Y$ is a $G$-equivariant isomorphism, where the action on the left is only on the first component. Such actions are called principal fiber bundles. As we consider only group schemes $G$ coming from a usual group (so that the corresponding group scheme is reduced), the morphism $Y\rightarrow X$ in the free case (and if everything is essentially of finite type) is also étale.
@@
where the lower row represents the setting of faithfully flat descent, see [41, Section VIII]. When the action is free, we have vertical isomorphisms everywhere, see [6, Example 6.2.B], and then the data above $Y$ (a module, an algebra, an affine scheme) with a group action are compatible with the group action if and only if they are descent data.
@@ 5. Invariant modules and invariant algebras
We have $S\otimes_{R}S\cong K^{|G|}\otimes S\cong S^{|G|}$ by [35, Proposition 0.9] and $R\rightarrow S$ is faithfully flat. We look at the diagram
@@
We now use descent properties for faithfully flat base change. By [26, Proposition 2.7.1 (viii)], it follows that the original ring homomorphism is an isomorphism, and by [26, Proposition 2.5.2] it follows that $M$ is locally free. ∎
@@
The previous result can also be deduced from faithfully flat descent for affine schemes (see [41, Théorème VIII.2.1]). To see this, one needs the mentioned result from GIT that $R\subseteq S$ is flat and that the compatible free group action corresponds to descent data, see [6, Example 6.2.B].
@@
Corollary 5.15.
Let $G$ be a finite group acting linearly and faithfully via $\beta$ on ${\mathbb{A}}^{d}$ with invariant ring $R=K[X_{1},\ldots,X_{d}]^{\beta}$. Let $V\subseteq{\mathbb{A}}^{d}$ be the nonempty open locus where the action $\beta$ is free and let $U\subseteq X$ be its image. Then, $(\operatorname{Spec}(\operatorname{Sym}_{R}(\Omega_{R|K})))|_{U}\cong(({\mathbb{A}}^{d}\times{\mathbb{A}}^{d})/\beta\times\beta))|_{U}$ is the tangent bundle over $U$.
Proof.
The linear action $\beta$ of the finite group $G$ on ${\mathbb{A}}^{d}$ induces the natural action of $G$ on the module of Kähler differentials $\Omega_{K[X_{1},\ldots,X_{d}]|K}$ and on the tangent bundle
$T_{{\mathbb{A}}^{d}}={\mathbb{A}}^{d}\times{\mathbb{A}}^{d}=\operatorname{Spec}(\operatorname{Sym}\Omega_{K[X_{1},\ldots,X_{d}]|K})\,$
which is given by $\beta\times\beta$. The corresponding invariant module is $\Omega_{R|K}$. The spectrum of its symmetric algebra is related to the quotient ${\mathbb{A}}^{d}\times{\mathbb{A}}^{d}/\beta\times\beta$ as described in Corollary 5.14. ∎
${\mathbb{A}}^{d}\times{\mathbb{A}}^{d}/\beta\times\beta$ is a natural candidate for the schematic closure of the tangent bundle of the regular locus of ${\mathbb{A}}^{d}/\beta$. However, in the case of a reflection group, this differs from the (trivial) tangent bundle of the quotient.
@@ 7. Ramification
Let $H\neq G$ be the kernel of the representation $\rho$. We mod out everywhere $H$, which gives the same situation back, with the extra property that $G/H$ acts faithfully on ${\mathbb{A}}^{m}$. Note that $Y/H$ is also nonreduced, otherwise, it would already be $\operatorname{Spec}K$. So we may assume that $G$ acts faithfully on ${\mathbb{A}}^{m}$. Let $V\subseteq{\mathbb{A}}^{m}$ be a nonempty open subset where the action is free, which exists by Lemma 2.3. We consider the action of $G$ on $Y\times V$. This action is also free by Lemma 2.4, so $Y\times V\rightarrow(Y\times V)/G$ is étale by Remark 2.2. Since $Y\times V$ is not reduced, $(Y\times V)/G$ is also not reduced by étale descent, see [46, Lemma 10.163.7], hence, $(Y\times{\mathbb{A}}^{m})/G$ is not reduced because it contains $(Y\times V)/G$ as an open subset. ∎
@@ 9. Regular representation
where the index set runs through a system $C$ of all irreducible representations, see [44, Corollary 1 in I.2.4]. $G$ acts on the polynomial ring $S[K^{|G|}]=\operatorname{Sym}(S^{|G|})$ with invariant algebra $(S[K^{|G|}])^{G}$. Here, $g$ sends the variable $W_{h}$ to $W_{hg}$.
@@ 10. Module schemes
Let $X$ denote a scheme. A well-known correspondence is that between (geometric) vector bundles over $X$ and locally free ${\mathcal{O}}_{X}$-modules on $X$. A vector bundle is a scheme $Z\rightarrow X$ that looks locally like $X\times{\mathbb{A}}^{r}\rightarrow X$ and has the property that the transition mappings are linear. A locally free ${\mathcal{O}}_{X}$-module looks locally like ${\mathcal{O}}_{X}^{r}$. The correspondence assigns to $Z$ the dual sheaf of the sheaf of sections and to a locally free sheaf $\mathcal{F}$ the (relative) spectrum of the symmetric algebra $\operatorname{Sym}({\mathcal{F}})=\oplus_{n\in{\mathbb{N}}}\operatorname{Sym}^{n}({\mathcal{F}})$, [28, Exercise II.5.18]. This last definition can be applied for any quasicoherent ${\mathcal{O}}_{X}$-module $\mathcal{F}$, see [23, Definition 1.7.8], and [23, Section 1.3] for the relative spectrum of a quasicoherent algebra.
@@
This formulation requires that we consider ${\mathbb{A}}^{1}_{X}$ with the natural addition and multiplication as a ring scheme (this term appears also in [36, Section 26], the affine line in the only ring scheme we are dealing with). If $\alpha:Z\times_{X}Z\rightarrow Z$ denotes the addition of the group scheme, then this definition means, to give some examples, that the diagram
@@
The concept of a module scheme is referred to briefly in [24, Chapitre 0.8] (in particular 0.8.2.3), Grothendieck writes “L’exemple des groupes a été traité avec assez de détails, mais par la suite nous laisserons généralement au lecteur le soin de développer des considérations analogues dans les exemples de structures algébriques que nous rencontrerons” (see [24, 8.2.8]). The concept is also mentioned in [14, 4.3.3], but a thorough study is missing. Related work, but only over a field, has been done by Greenberg ([21], [22], where module variety over a (noncommutative) ring variety is used) and, in the context of graded Hop algebras, by Milnor and Moore ([34], see also Remark 11.4 below). See also [40] for the notion of schéma en $F$-vectoriels.
The following lemma is stated in [23, 1.7.13] (see also [27, Section 9.4]. Grothendieck sets $V({\mathcal{E}})=\operatorname{Spec}(\operatorname{Sym}{\mathcal{E}})$ and calls this, quite inappropriately, fibré vectoriel ([14] uses fibration vectorielle instead), and writes “Nous interpréterons ces faits en disant que $S[T]$ est un $S$-schéma d’anneaux et que $V({\mathcal{E}})$ est un $S$-schéma de modules sur le S-schéma d’anneaux (cf. chap. 0, § 8)”.
@@ 11. Module schemes and graded Hopf algebras
Positively graded Hopf-algebras $A$ with $A_{0}=K$ a field (called connected) appear in the work of Milnor and Moore. If $K$ has characteristic 00, they prove results (the theorem of Milnor-Moore) that imply (under commutative and cocommutative assumption) that $A$ is a polynomial algebra over $K$, generated by homogeneous elements, see [34, Corollary 4.18] and [12, Theorem 3.8.3]. Translated into Grothendiecks terminology, this means via Lemma 11.1 and Lemma 11.3 that a module scheme over the spectrum of a field of characteristic zero is a vector space (the spectrum of the symmetric algebra of a vector space).
@@ 12. Module schemes up to normalization
The data that constitute a module scheme up to normalization are determined by the data on the restriction to $U$, by [27, Proposition 7.2.2]. The only relevant question is whether these data can be extended to some $\widetilde{Z\times_{X}Z}$. Since the properties are expressed by the commutativity of certain diagrams, i.e., identities between morphisms, these identities only need to be checked on a dense open subset.
@@ 13. Linear actions on module schemes
(1) is clear. (2) and (3) follow from Lemma 12.5. (4) is an application of [27, Proposition 7.2.2]. For $U$ sufficiently small, we get module schemes. ∎
@@
A symmetric algebra $\operatorname{Spec}\operatorname{Sym}M\rightarrow\operatorname{Spec}R$ for a finitely generated $R$-module $M$ comes naturally together with a linear injective mapping $\operatorname{Spec}\operatorname{Sym}M\rightarrow\operatorname{Spec}R\times{\mathbb{A}}^{\nu}$ above $\operatorname{Spec}R$, defined by a surjection $R^{\nu}\rightarrow M$, where $\nu$ is the number of generators of the module. If $M$ is locally free above $U$, then we get, by [27, Théorem 9.7.4], on $U$ a mapping $\vartheta:U\rightarrow\operatorname{Gr}(\nu+r,r)$ into the Grassmann variety, where $r$ denotes the rank of $M$, and where the pull-back of the universal bundle on $\operatorname{Gr}(\nu+r,r)$ gives back the embedding of the bundle over $U$. The higherdimensional fibers of $\operatorname{Spec}\operatorname{Sym}M\rightarrow\operatorname{Spec}R$ outside of $U$ are reflected by the property that the closure of the image of $\vartheta$ degenerates to a higher dimension. For a module scheme $\operatorname{Spec}A\rightarrow\operatorname{Spec}R$ up to normalization, realizing the same module $M$ above $U$, the embedding $\operatorname{Spec}A\subseteq\operatorname{Spec}R\times{\mathbb{A}}^{\mu}$ (where $\mu$ denotes the number of $R$-algebra generators of $A$) is not linear in general (not even above $U$), as Example 3.10 shows. Hence, there is also no interpretation of $\operatorname{Spec}A$ using the Grassmannian.
@@ 16. Similar modifications
In the construction of abelian varieties by A. Weil, one first has only a rational composition $V\times V\cdots\rightarrow V$, which must then be extended to a proper variety (see [51, Théorème 15], [14, Exposé XVIII]), and a similar construction occurs if one wants to spread an abelian variety defined over a quotient field $Q(R)$ to an abelian scheme over a $R$, see [6, Section 4.3].